OpenStudy (anonymous):

find an angle such that 3 times the complement of the angle is 50 degrees greater than the supplement of the angle find an angle such that 3 times the complement of the angle is 50 degrees greater than the supplement of the angle @Mathematics

OpenStudy (anonymous):

I'm not positive, but I think you can set it up like this: Angle=x 3(90-x)-50=180-x Then solve for x

OpenStudy (anonymous):

270-3x-50=180-x 40=2x x=20

OpenStudy (anonymous):

let the angle be X 3(90-X) = (180-X) +50 solving X = 20